Binding Company reported net fixed assets of $190,000 at the end of Year 1 and $210,000 at the end of Year 2. The company also reported net revenues of $200,000 for Year 1 and $240,000 for Year 2. What was the fixed asset turnover ratio? (Round your answer to 1 decimal point.)
Answer #1

Fixed asset turnover ratio = Net revenues / Average fixed assets

Year 1 = 200,000 / 190,000 = 1.1 times

Year 2 = 240,000 / [ ( 190,000 + 210,000)/2] = 1.2 times